OpenStudy (anonymous):

Determine whether parallelogram JKLM with vertices J(-1, -1), K(4, 4), L(9, -1) and M(4, -6) is a rhombus, square, rectangle or all three.

OpenStudy (anonymous):

You part of keystone?

OpenStudy (jdoe0001):

what do you think?

OpenStudy (kira_yamato):

 \[AB = \sqrt{\left( x _{B} - x_{A} \right)^2 + \left( y _{B} - y_{A} \right)^2}\] Use this formula to find JK and KL, and determine if they are of the same length

OpenStudy (anonymous):

I ended up with JK = 5√2 Still working on KL

OpenStudy (anonymous):

They're equal

OpenStudy (anonymous):

It's all three, check your message box for why.

OpenStudy (kira_yamato):

Sorry it should be JK vector dot LK vector, but you'll get the same answer anyway